# Network-Optimised Spiking Neural Network for Event-Driven Networking

### v3

@@ Appendix I Stochastic arrivals: shot noise model and calibration
Figure 14: Equilibrium $v^{*}$ versus input $I$ with saddle–node ($\blacksquare$) and Hopf ($\blacktriangle$) markers. (Left) Increasing service rate $\lambda$ shifts both onsets to higher II. (Middle) Increasing excitability $\alpha$ advances both onsets. (Right) Varying recovery coupling $b$ changes the balance between abab and $a+\mu$: when $ab>(a+\mu)^{2}$ an oscillatory Hopf branch is admissible, otherwise only an SN appears. Stable branches are shown in colour; unstable in faint grey. The exact marker coordinates used here are reported in Table 7.
@@
$\inf_{v\in\mathcal{I}}\bigl(f^{\prime}_{\mathrm{sat}}(v)+L\bigr)$ $>0,$ (87)
then $F$ is strictly increasing on $\mathcal{I}$ and has at most one root. If, in addition, there exist $v_{\pm}\in\mathcal{I}$ with $F(v_{-})\,F(v_{+})\leq 0$, then a unique equilibrium $v^{*}\in\mathcal{I}$ exists. A sufficient global condition is
$L$ $>-\,\frac{3\sqrt{3}}{8}\,\frac{\alpha}{\sqrt{\kappa}}.$ (88)
@@
so that $\tilde{v}$ and $\tilde{t}$ are dimensionless queue level and time. This scaling lets parameters fitted on sites with different sampling and buffer sizes be compared on a common basis and leaves the network-level stability proxy $g\,\rho(W)\,f^{\prime}_{\mathrm{sat}}(v^{*})<\Lambda$ invariant. Choose $V$ as the full-buffer level used in the queue normalisation (for example, bytes or packets mapped to $v\in[0,1]$), and choose $T$ as the dominant local timescale: the scheduler epoch, the service drain time $1/\lambda$, or the telemetry bin width when that is the binding constraint. With (92), the left-hand side of (2) satisfies
@@
$f_{\mathrm{sat}}(v)\;=\;\frac{\alpha\,v^{2}}{1+\kappa v^{2}}\;\rightarrow\;\tilde{f}_{\mathrm{sat}}(\tilde{v})\;=\;\frac{T}{V}\,f_{\mathrm{sat}}(V\tilde{v})\;=\;\frac{\tilde{\alpha}\,\tilde{v}^{2}}{\,1+\tilde{\kappa}\,\tilde{v}^{2}\,},\quad\tilde{\alpha}\;=\;\alpha TV,\;\;\tilde{\kappa}\;=\;\kappa V^{2}.$ (94)
@@
$\tilde{\beta}\;=\;\beta T,\quad\tilde{\lambda}\;=\;\lambda T,\quad\tilde{\chi}\;=\;\chi T,\;\;\tilde{\gamma}\;=\;\frac{\gamma T}{V},\quad\tilde{v}_{\mathrm{rest}}\;=\;\frac{v_{\mathrm{rest}}}{V}.$ (95)
@@
$\frac{du}{dt}\;=\;a(bv-u)-\mu u\;\rightarrow\;\frac{d\tilde{u}}{d\tilde{t}}\;=\;\tilde{a}\bigl(\tilde{b}\,\tilde{v}-\tilde{u}\bigr)-\tilde{\mu}\,\tilde{u},\;\tilde{a}\;=\;aT,\;\;\tilde{\mu}\;=\;\mu T,\;\tilde{b}\;=\;b,\;\;\tilde{u}\;=\;\frac{u}{V}$ (96)
@@
$\tilde{I}_{i}(\tilde{t})\;=\;\frac{T}{V}\,I_{i}(t),\qquad\tilde{\tau}_{ij}\;=\;\frac{\tau_{ij}}{T},\;\;\tilde{w}_{ij}\;=\;w_{ij}\;\;\text{(dimensionless by construction)}.$ (97)
@@
$\tilde{\eta}_{i}(\tilde{t})=\frac{T}{V}\,\eta_{i}(t)\;=\;\sum_{n=1}^{N_{i}(t)}\tilde{A}_{i,n}\,\tilde{\kappa}_{s}\bigl(\tilde{t}-\tilde{t}_{i,n}\bigr),\;\;\tilde{A}_{i,n}\;=\;\frac{T}{V}\,A_{i,n},\;\;\tilde{\kappa}_{s}(\tilde{t})\;=\;e^{-\tilde{t}/\tilde{\tau}_{s}}H(\tilde{t}),\;\;\tilde{\tau}_{s}\;=\;\frac{\tau_{s}}{T}.$ (98)
@@
The local slope bound in (34) and the service–damping aggregate in (90) feed directly into the small-signal analysis in K.2 and the network coupling results in §3.3. Let $\rho(W)$ be the spectral radius of the coupling matrix and $g$ any global gain applied to $W$. Linearisation of (2)–(3) about an operating point $v^{*}$ yields a block Jacobian whose dominant network term is proportional to $g\,\rho(W)\,f^{\prime}_{\mathrm{sat}}(v^{*})$. Compare this against the net drain
@@
$g\,\rho(W)\,f^{\prime}_{\mathrm{sat}}(v^{*})$ $<\Lambda.$ (100)
Under the rescaling (92), both sides scale by the same factor: $T\,f^{\prime}_{\mathrm{sat}}$ and $T\,\Lambda$ are dimensionless, so (100) is invariant and can be checked in either set of units. The proxy exposes the same levers used in operations. One can reduce $\rho(W)$ by reweighting or sparsifying couplings, reduce $g$ through gain scheduling, or raise $\Lambda$ by increasing service $\lambda$, damping $\chi$, or the recovery rate $a$ (with the trade-off set by $b$ and $\mu$). It is useful to track the operational margin
$\Delta_{\mathrm{net}}$ $:=\Lambda-g\,\rho(W)\,f^{\prime}_{\mathrm{sat}}(v^{*}),$ (101)
@@
Finite networks smooth the bifurcation transitions seen in the scalar reduction. We quantify collective timing with a Kuramoto-type order parameter
$R(t)\;=\;\frac{1}{N}\Bigl|\sum_{j=1}^{N}e^{\mathrm{i}\phi_{j}(t)}\Bigr|,$
where spike phases $\phi_{j}(t)$ are obtained by linear phase interpolation between successive spikes of node $j$ (phase resets to 0 at a spike, advances to $2\pi$ at the next). Figure 15 plots the time average $\langle R\rangle$ against the effective coupling $k=g\,\rho(W)$ for several network sizes $N$. The curves show the expected S-shaped rise: for small $k$ the system is desynchronised and $\langle R\rangle$ is low; once $k$ passes a critical value the order parameter increases rapidly and then saturates.
@@
10:  (3) State update (forward Euler): for each node $i$
$v_{i}^{\text{old}}\leftarrow v_{i},\qquad u_{i}^{\text{old}}\leftarrow u_{i}$
$v_{i}\leftarrow v_{i}^{\text{old}}+dt\cdot\Bigl(f_{\mathrm{sat}}(v_{i}^{\text{old}})+\beta v_{i}^{\text{old}}+\gamma-u_{i}^{\text{old}}+I_{i}(t)-\lambda v_{i}^{\text{old}}-\chi\bigl(v_{i}^{\text{old}}-v_{\mathrm{rest}}\bigr)\Bigr)$
$u_{i}\leftarrow u_{i}^{\text{old}}+dt\cdot\Bigl(a\bigl(bv_{i}^{\text{old}}-u_{i}^{\text{old}}\bigr)-\mu u_{i}^{\text{old}}\Bigr)$
@@
12:  (4) Threshold and reset: for each node $i$
$v_{\mathrm{th},i}(t)\leftarrow v_{\mathrm{th}}+\sigma\,\xi_{i}(t)\qquad(\xi_{i}(t)\sim\mathcal{N}(0,1),\ \sigma=0\text{ for deterministic runs})$
@@
18:   Soft exponential reset:
$v_{i}\leftarrow c+(v_{i}-c)\,e^{-r_{\mathrm{reset}}dt},\qquad u_{i}\leftarrow u_{i}+d$

### v4

@@ Appendix I Stochastic arrivals: shot noise model and calibration
Figure 14: Equilibrium $v^{*}$ versus input II with saddle–node ($\blacksquare$) and Hopf ($\blacktriangle$) markers. (Left) Increasing service rate $\lambda$ shifts both onsets to higher II. (Middle) Increasing excitability $\alpha$ advances both onsets. (Right) Varying recovery coupling $b$ changes the balance between abab and $a+\mu$: when $ab>(a+\mu)^{2}$ an oscillatory Hopf branch is admissible, otherwise only an SN appears. Stable branches are shown in colour; unstable in faint grey. The exact marker coordinates used here are reported in Table 7.
@@
$\sup_{v\in\mathcal{I}}\bigl(f^{\prime}_{\mathrm{sat}}(v)+L\bigr)$ $<0,$ (102)
then $F$ is strictly decreasing on $\mathcal{I}$ and has at most one root. If, in addition, there exist $v_{-},v_{+}\in\mathcal{I}$ with $F(v_{-})\,F(v_{+})\leq 0$, then a unique equilibrium $v^{*}\in\mathcal{I}$ exists.
A sufficient global condition for (102) (using $\sup_{v}f^{\prime}_{\mathrm{sat}}(v)\leq\frac{3\sqrt{3}}{8}\frac{\alpha}{\sqrt{\kappa}}$ for $\kappa>0$) is
$L$ $<-\,\frac{3\sqrt{3}}{8}\,\frac{\alpha}{\sqrt{\kappa}}.$ (103)
@@
so that $\tilde{v}$ and $\tilde{t}$ are dimensionless queue level and time. This scaling lets parameters fitted on sites with different sampling and buffer sizes be compared on a common basis and leaves the network-level stability proxy $g\,\rho(W)\,+f^{\prime}_{\mathrm{sat}}(v^{*})<\Lambda$ invariant. Choose $V$ as the full-buffer level used in the queue normalisation (for example, bytes or packets mapped to $v\in[0,1]$), and choose $T$ as the dominant local timescale: the scheduler epoch, the service drain time $1/\lambda$, or the telemetry bin width when that is the binding constraint. With (107), the left-hand side of (2) satisfies
@@
$\begin{split}f_{\mathrm{sat}}(v)\;=\;\frac{\alpha\,v^{2}}{1+\kappa v^{2}}\;\rightarrow\;\tilde{f}_{\mathrm{sat}}(\tilde{v})\\ \;=\;\frac{T}{V}\,f_{\mathrm{sat}}(V\tilde{v})\;=\;\frac{\tilde{\alpha}\,\tilde{v}^{2}}{\,1+\tilde{\kappa}\,\tilde{v}^{2}\,},\\ \tilde{\alpha}\;=\;\alpha TV,\;\;\tilde{\kappa}\;=\;\kappa V^{2}.\end{split}$ (109)
@@
$\begin{split}\tilde{\beta}\;=\;\beta T,\quad\tilde{\lambda}\;=\;\lambda T,\quad\tilde{\chi}\;=\;\chi T,\\ \tilde{\gamma}\;=\;\frac{\gamma T}{V},\quad\tilde{v}_{\mathrm{rest}}\;=\;\frac{v_{\mathrm{rest}}}{V}.\end{split}$ (110)
@@
$\begin{split}\frac{du}{dt}\;=\;a(bv-u)-\mu u\;\rightarrow\;\frac{d\tilde{u}}{d\tilde{t}}\;=\;\tilde{a}\bigl(\tilde{b}\,\tilde{v}-\tilde{u}\bigr)\\ -\tilde{\mu}\,\tilde{u},\;\tilde{a}\;=\;aT,\;\;\tilde{\mu}\;=\;\mu T,\;\tilde{b}\;=\;b,\;\;\tilde{u}\;=\;\frac{u}{V}\end{split}$ (111)
@@
$\begin{split}\tilde{I}_{i}(\tilde{t})\;=\;\frac{T}{V}\,I_{i}(t),\;\;\;\;\tilde{\tau}_{ij}\;=\;\frac{\tau_{ij}}{T},\\ \tilde{w}_{ij}\;=\;w_{ij}\;\;\text{(dimensionless by construction)}.\end{split}$ (112)
@@
$\begin{split}\tilde{\eta}_{i}(\tilde{t})=\frac{T}{V}\,\eta_{i}(t)\;=\;\sum_{n=1}^{N_{i}(t)}\tilde{A}_{i,n}\,\tilde{\kappa}_{s}\bigl(\tilde{t}-\tilde{t}_{i,n}\bigr),\\ \tilde{A}_{i,n}\;=\;\frac{T}{V}\,A_{i,n},\;\;\tilde{\kappa}_{s}(\tilde{t})\;=\;e^{-\tilde{t}/\tilde{\tau}_{s}}H(\tilde{t}),\;\;\tilde{\tau}_{s}\;=\;\frac{\tau_{s}}{T}.\end{split}$ (113)
@@
The local slope bound in (33) and the service–damping aggregate in (105) feed directly into the small-signal analysis in K.3 and the network coupling results in §3.3. Let $\rho(W)$ be the spectral radius of the coupling matrix and $g$ any global gain applied to $W$. Linearisation of (2)–(3) about an operating point $v^{*}$ yields a block Jacobian whose dominant network term is proportional to $g\,\rho(W)\,+f^{\prime}_{\mathrm{sat}}(v^{*})$. Compare this against the net drain
@@
$g\,\rho(W)$ $<\Lambda-f^{\prime}_{\mathrm{sat}}(v^{*}),$ (115)
equivalently $g\,\rho(W)+f^{\prime}_{\mathrm{sat}}(v^{*})<\Lambda$. Under the rescaling (107), both $T\,f^{\prime}_{\mathrm{sat}}$ and $T\,\Lambda$ are dimensionless, so (115) is invariant and can be checked in either set of units. The proxy exposes the same levers used in operations. One can reduce $\rho(W)$ by reweighting or sparsifying couplings, reduce $g$ through gain scheduling, or raise $\Lambda$ by increasing service $\lambda$, damping $\chi$, or the recovery rate $a$ (with the trade-off set by $b$ and $\mu$). It is useful to track the operational margin
$\Delta_{\mathrm{net}}$ $:=\Lambda-f^{\prime}_{\mathrm{sat}}(v^{*})-g\,\rho(W).$ (116)
@@
Finite networks smooth the bifurcation transitions seen in the scalar reduction. We quantify collective timing with a Kuramoto-type order parameter $R(t)\;=\;\frac{1}{N}\Bigl|\sum_{j=1}^{N}e^{\mathrm{i}\phi_{j}(t)}\Bigr|,$ where spike phases $\phi_{j}(t)$ are obtained by linear phase interpolation between successive spikes of node $j$ (phase resets to 0 at a spike, advances to $2\pi$ at the next). Figure 15 plots the time average $\langle R\rangle$ against the effective coupling $k=g\,\rho(W)$ for several network sizes $N$. The curves show the expected S-shaped rise: for small $k$ the system is desynchronised and $\langle R\rangle$ is low; once $k$ passes a critical value the order parameter increases rapidly and then saturates.
@@
10:  (3) State update (forward Euler): for each node $i$ $v_{i}^{\text{old}}\leftarrow v_{i},\qquad u_{i}^{\text{old}}\leftarrow u_{i}$ $v_{i}\leftarrow v_{i}^{\text{old}}+dt\cdot\Bigl(f_{\mathrm{sat}}(v_{i}^{\text{old}})+\beta v_{i}^{\text{old}}+\gamma-u_{i}^{\text{old}}+I_{i}(t)-\lambda v_{i}^{\text{old}}-\chi\bigl(v_{i}^{\text{old}}-v_{\mathrm{rest}}\bigr)\Bigr)$ $u_{i}\leftarrow u_{i}^{\text{old}}+dt\cdot\Bigl(a\bigl(bv_{i}^{\text{old}}-u_{i}^{\text{old}}\bigr)-\mu u_{i}^{\text{old}}\Bigr)$
@@
12:  (4) Threshold and reset: for each node $i$ $v_{\mathrm{th},i}(t)\leftarrow v_{\mathrm{th}}+\sigma\,\xi_{i}(t)\qquad(\xi_{i}(t)\sim\mathcal{N}(0,1),\ \sigma=0\text{ for deterministic runs})$
@@
18:   Soft exponential reset: $v_{i}\leftarrow c+(v_{i}-c)\,e^{-r_{\mathrm{reset}}dt},\qquad u_{i}\leftarrow u_{i}+d$